Derive the equation of

 A)Q(t) for discharging process

B)V(t) for discharging process

C)I(t) for discharging process

D)show discharging graph

E)what happen for Q(t) for t=goes to +ve infinity for discharging process.


1
Expert's answer
2021-07-08T09:57:03-0400

1)


"Q=Q_0e^{-\frac{t}{RC}}"

2)


"V(t)=\frac{Q_0}{C}e^{-\frac{t}{RC}}"

3)


"|i(t)|=\frac{Q_0}{RC}e^{-\frac{t}{RC}}"

4)



5) It will be zero.